A jet plane travelling 1890 km/hr (525 m/s) pulls out of a dive by moving in an arc of radius 5.20 km. What is the plane’s acceleration in g’s?
1
Expert's answer
2019-01-21T11:17:09-0500

The centripetal acceleration

a=v^2/R

So

a=(525 m/s)^2/(5 200 m)=53 m/s^2 =5.4g